# Communication Constellation Design of Minimum Number of Satellites with Continuous Coverage and Inter-Satellite Link

### v2

@@ 2.1 Walker Constellation
From the equation (1), the RAAN is equally spaced by $\Delta\Omega=360/P\deg$ in the range $\Omega_{m}\in\left[0,360\right]$ $\deg$. The intraplane spacing $\Delta M=360/S\deg$ generates a symmetric location in the orbital plane in the range $M\in\left[0,360\right]\deg$ and is equivalent to the relative angular distance between the satellites in the orbital plane. The relative argument of the latitude ($\Delta u$) in the adjacent orbital planes is derived as $\Delta u=360/T\cdot F\deg$ from Eq. (2).
@@ 2.1.1 Pattern Repetition Period
Let us define the set of satellites in the $m$th orbital plane as $\textbf{N}=\left\{n\mid n=1,2,...,S\right\}$. Then, Equations (10) and (11) derive
@@ 2.1.2 Duplicate Allocation of Orbital Elements
$\begin{cases}m^{\prime}=1+\frac{P}{2}\\ n^{\prime}=1+\frac{S-F}{2}.\end{cases}$ (14)
Therefore, the Walker-Delta design parameters that accommodate Eq. (14) must be avoided in the design procedure.
@@ 5.2.2 Common Ground-track Constellation
Figure 10: (a) Minimum and (b) maximum relative distance between the adjacent orbital plane of quasi-symmetric and BILP CGT constellations
The relative motion equations (48), (50), and (51) imply that the relative motion of two orbits with the same altitude and inclination is a function of the inclination, relative RAAN, and relative argument of latitude. Let us define a set of temporal location $\tau_{k}$ in Eq. (26) as $\boldsymbol{\tau}_{k}$. Then, $\boldsymbol{\tau}_{k,qs}$ and $\boldsymbol{\tau}_{k,bilp}$ are calculated as Eqs. (53) and (54). The temporal location $\tau_{k}$ derives the RAAN $\Omega_{k}$ as shown in Eq. (55)[7].
@@
where the subscript ‘0’ means that the variable is relevant to the seed satellite. The equation (25) describes the relationship between $\Omega_{k}$ and $M_{k}$.
@@
Thus, the equations (53), (54), and (56) are used to calculate $\Delta\tau_{k}$ for the two CGT constellations as
@@
The equation (55) derives the relative RAAN $\Delta\Omega_{k}$ and the relative mean anomaly $\Delta M_{k}$ as
@@
The minimum and maximum relative distances in Eqs. (46) and (47) are depicted in Figure 10. The left and right panels appear almost identical because the difference between the minimum and maximum distances is less than 1000 km. The minimum and maximum distances for the quasi-symmetric CGT constellation are 13854.32 and 13886.49 km, respectively. The BILP CGT constellation has various values of $\Delta\tau_{k}$. When $\Delta\tau_{k}$ is 8, the minimum and maximum distances are 13164.56 and 13191.33 km.
@@ 6 Conclusion
This paper investigates the continuous coverage analysis methods and inter-satellite link connectivity analysis methods for communication satellite constellations. The bounded Voronoi diagram designs a homogeneous constellation to ensure continuous regional and global coverage. The APC decomposition can be implemented for the CGT constellation’s coverage analysis. The relative motion in adjacent orbital planes derives the analytical solutions that should be constrained within the inter-satellite link range. The Walker-Delta constellation and two types of CGT constellations are applied as examples. As a result, the Walker-Delta constellation design method consists of more satellites than CGT constellations for a single-fold coverage. However, the relative motion range is shorter and more consistent, which implies that the Walker-Delta constellation has advantages for inter-satellite links. The quasi-symmetric and BILP methods have been suggested for the CGT constellation design method. The BILP optimal CGT constellation has an asymmetric configuration but achieves a smaller number of satellites. The relative motion of BILP constellation has a variety of ranges due to its asymmetricity, but satellites are located closer than the quasi-symmetric constellation. In summary, the BILP constellation is advantageous in terms of the number of satellites required for a single-fold coverage. However, the Walker-Delta constellation may have a shorter and more rigorous relative motion range, which is beneficial for inter-satellite links.

### v3

@@ 2.1 Walker Constellation
From the Eq. (1), the RAAN is equally spaced by $\Delta\Omega=360/P\deg$ in the range $\Omega_{m}\in\left[0,360\right]$ $\deg$. The intraplane spacing $\Delta M=360/S\deg$ generates a symmetric location in the orbital plane in the range $M\in\left[0,360\right]\deg$ and is equivalent to the relative angular distance between the satellites in the orbital plane. The relative argument of the latitude ($\Delta u$) in the adjacent orbital planes is derived as $\Delta u=360/T\cdot F\deg$ from Eq. (2).
@@ 2.1.1 Pattern Repetition Period
Let us define the set of satellites in the $m$th orbital plane as $\textbf{N}=\left\{n\mid n=1,2,...,S\right\}$. Then, Eqs. (10) and (11) derive
@@ 2.1.2 Duplicate Allocation of Orbital Elements
$\begin{cases}m^{\prime}=1+\frac{P}{2}\\ n^{\prime}=mod\left(1+\frac{S-F}{2},S\right)+S\cdot\delta\left(mod\left(1+\frac{S-F}{2},S\right),0\right).\end{cases}$ (14)
where $mod\left(x,y\right)$ denotes the modulo operation, which returns the remainder when $x$ is divided by $y$, and $\delta\left(x,y\right)$ represents the Kronecker delta function, which equals $1$ if $x=y$ and 00 otherwise. Therefore, the Walker-Delta design parameters that accommodate Eq. (14) must be avoided in the design procedure.
@@ 5.2.2 Common Ground-track Constellation
Figure 10: Minimum and maximum relative distance between the adjacent orbital plane of quasi-symmetric and BILP CGT constellations
The relative motion equations, Eqs. (48), (50), and (51), imply that the relative motion of two orbits with the same altitude and inclination is a function of the inclination, relative RAAN, and relative argument of latitude. Let us define a set of temporal location $\tau_{k}$ in Eq. (26) as $\boldsymbol{\tau}_{k}$. Then, $\boldsymbol{\tau}_{k,qs}$ and $\boldsymbol{\tau}_{k,bilp}$ are calculated as Eqs. (53) and (54). The temporal location $\tau_{k}$ derives the RAAN $\Omega_{k}$ as shown in Eq. (55)[7].
@@
where the subscript ‘0’ means that the variable is relevant to the seed satellite. Equation (25) describes the relationship between $\Omega_{k}$ and $M_{k}$.
@@
Thus, Eqs. (53), (54), and (56) are used to calculate $\Delta\tau_{k}$ for the two CGT constellations as
@@
Equation (55) derives the relative RAAN $\Delta\Omega_{k}$ and the relative mean anomaly $\Delta M_{k}$ as
@@
The minimum and maximum relative distances in Eqs. (46) and (47) are depicted in Figure 10. The minimum and maximum relative distances appear almost identical for some cases because the differences are less than 1000 km. The minimum and maximum relative distances for the quasi-symmetric CGT constellation are 13854.32 and 13886.49 km, respectively. The BILP CGT constellation has various values of $\Delta\tau_{k}$. When $\Delta\tau_{k}$ is 8, the minimum and maximum relative distances reach their largest values at 13164.56 and 13191.33 km, respectively.
@@ 6 Conclusion
This paper investigates the continuous coverage analysis methods and the inter-satellite link connectivity analysis method for communication satellite constellations. The bounded Voronoi diagram is used to design a homogeneous constellation that ensures continuous regional and global coverage. The APC decomposition, based on the grid method, can be implemented for the CGT constellation’s coverage analysis. The relative motion in adjacent orbital planes yields the analytical solutions that must be constrained within the inter-satellite link range.
The coverage performance of the Walker-Delta constellation was analyzed using the bounded Voronoi diagram as an example. Two types of CGT constellations, the quasi-symmetric and the BILP, were analyzed using APC decomposition to compare their coverage performance. The BILP optimal CGT constellation has an asymmetric configuration but achieves a smaller number of satellites. However, the relative motion range of the Walker-Delta constellation is shorter and more consistent, which implies that the Walker-Delta constellation has advantages for inter-satellite links. The relative motion of the BILP constellation has a variety of ranges due to its asymmetry, but satellites are located closer than the quasi-symmetric constellation. In summary, the BILP constellation is advantageous in terms of the number of satellites required for a single-fold coverage. However, the Walker-Delta constellation may have a shorter and more stable relative motion range, which is beneficial for inter-satellite links.